solve (1+t²)y' +4ty=(1+t²)^-2; y(0)=1


1
Expert's answer
2022-03-20T06:43:43-0400

"\text{ Solve the linear equation } \frac{d y(t)}{d t}\left(t^{2}+1\right)+4 t y(t)=\frac{1}{\left(t^{2}+1\right)^{2}}, \text{ such that } y(0)=1 :\\[4mm]
\text{ Divide both sides by } t^{2}+1 :\\[1.5mm]
\frac{d y(t)}{d t}+\frac{4 t y(t)}{t^{2}+1}=\frac{1}{\left(t^{2}+1\right)^{3}}\\[2mm]
\text{ Let } \mu(t)=e^{\int(4 t) /\left(t^{2}+1\right) d t}=\left(t^{2}+1\right)^{2}.\\[4mm]
\text{ Multiply both sides by } \mu(t) :\\[1.5mm]
\left(t^{2}+1\right)^{2} \frac{d y(t)}{d t}+\left(4 t\left(t^{2}+1\right)\right) y(t)=\frac{1}{t^{2}+1}\\[2mm]
\text{ Substitute } 4 t\left(t^{2}+1\right)=\frac{d}{d t}\left(\left(t^{2}+1\right)^{2}\right) :\\[2mm]
\left(t^{2}+1\right)^{2} \frac{d y(t)}{d t}+\frac{d}{d t}\left(\left(t^{2}+1\right)^{2}\right) y(t)=\frac{1}{t^{2}+1}\\[4mm]
\text{ Apply the reverse product rule } f \frac{d g}{d t}+g \frac{d f}{d t}=\frac{d}{d t}(f g) \text{ to the left-hand side:}\\[2mm]
\frac{d}{d t}\left(\left(t^{2}+1\right)^{2} y(t)\right)=\frac{1}{t^{2}+1}"


"\text{Integrate both sides with respect to } t :\\[1.5mm]
\int \frac{d}{d t}\left(\left(t^{2}+1\right)^{2} y(t)\right) d t=\int \frac{1}{t^{2}+1} d t\\[4mm]

\text{Evaluate the integrals:}\\
\left(t^{2}+1\right)^{2} y(t)=\tan ^{-1}(t)+c_{1}, \text{ where } c_{1} \text{ is an arbitrary constant.}\\[4mm]
\text{Divide both sides by } \mu(t)=\left(t^{2}+1\right)^{2} :\\[1.5mm]
y(t)=\frac{\tan ^{-1}(t)+c_{1}}{\left(t^{2}+1\right)^{2}}\\[4mm]

\text{ Solve for } c_{1} \text{ using the initial conditions:}\\[2mm]
\text{ Substitute } y(0)=1 \text{ into } y(t)=\frac{\tan ^{-1}(t)+c 1}{\left(t^{2}+1\right)^{2}} :\\
c_{1}=1\\[2mm]
\text{ Substitute } c_{1}=1 \text{ into } y(t)=\frac{\tan ^{-1}(t)+c_{1}}{\left(t^{2}+1\right)^{2}}\\[2mm]
\text{ Answer:}\\[2mm]
y(t)=\frac{\tan ^{-1}(t)+1}{\left(t^{2}+1\right)^{2}}"